# Recovering Pulsar Braking Index from a Population of Millisecond Pulsars

### v1

@@ Abstract
The braking index, $n$, of a pulsar is a measure of its angular momentum loss and the value it takes corresponds to different spin-down mechanisms. For a pulsar spinning down due to gravitational wave emission from the principal mass quadrupole mode alone, the braking index would equal exactly 5. Unfortunately, for millisecond pulsars, it can be hard to measure observationally due to the extremely small second time derivative of the rotation frequency, $\ddot{f}$. This paper aims to examine whether it could be possible to extract the distribution of $n$ for a whole population of pulsars rather than measuring the values individually. We use simulated data with an injected $n=5$ signal for 47 millisecond pulsars and extract the distribution using hierarchical Bayesian inference methods. We find that while possible, observation times of over 20 years and RMS noise of the order of $10^{-5}$ ms are needed, which can be compared to the mean noise value of $3\!\times\!10^{-4}$ ms for the recent wideband 12.5-year NANOGrav sample (Alam et al., 2020), which provided the pulsar timing data used in this paper.
facilities: NANOGrav (Alam et al., 2020, 2021)software: enterprise (Ellis et al., 2020), enterprise_extensions (Taylor et al., 2021, 2024), enterprise_warp (Goncharov, 2021; Goncharov et al., 2024) posteriorstacker (Buchner, 2021), bilby (Ashton et al., 2019; Romero-Shaw et al., 2020; Ashton & Talbot, 2021), numpy (Harris et al., 2020), Astropy (Robitaille et al., 2013; Price-Whelan et al., 2018), Tempo2 (Hobbs et al., 2006; Edwards et al., 2006; Hobbs et al., 2009), matplotlib (Hunter, 2007),
@@ 1 Introduction
$\dot{f}_{\rm dip}=-\frac{2}{3c^{3}}\frac{B^{2}}{I_{zz}}R^{6}f^{3}\sin{\alpha},$ (2)
which gives a braking index of 3 (e.g., Hamil et al., 2015). Here, $c$ is the speed of light, $B$ represents the magnitude of the dipolar magnetic field at the surface, $I_{zz}$ is the moment of inertia, $\alpha$ is the angle between the magnetic field axis and the rotation axis, and $R$ is the star’s radius.
@@
$\dot{f}_{\rm GW}=-\frac{32}{5}\frac{G}{c^{5}}I_{zz}\varepsilon^{2}f^{5},$ (3)
@@ 2.1 Creating TOAs
A sample of $47$ pulsars from the wideband NANOGrav 12.5-year data set (Alam et al., 2020, 2021) are used as example pulsars for this analysis. All pulsars in this sample are MSPs, meaning they have a rotation frequency of over 100 Hz. This puts the sample closer to the cut-off described in Woan et al. (2018) which is caused by $n=5$ processes than if it contained pulsars with lower frequencies. It may be noted that 30 of the pulsars are in binary systems; however, this information is not used in this study and they are treated as isolated pulsars.
@@
In this analysis, we allow our simulated TOAs to be created with a root mean square (RMS) noise value for white noise alone. For each pulsar, the noise value is taken from Table 5 in Alam et al. (2020). In cases where the source has a separate red noise value, the value for white noise is taken only. As the effect of timing noise on the accuracy of this method will be tested, we need to be able to vary the average noise of the sample to any desired value $\text{RMS}_{\text{desired}}$. Therefore, we scale the RMS noise of each individual pulsar using
@@
where RMSi is the RMS noise value for pulsar $i$ taken from Alam et al. (2020), $\mu$ is the mean of the original dataset, and $\mu_{\text{desired}}$ is the mean the sample will be scaled to. This retains the realistic natural variation of RMS noise throughout the population.
@@ 2.3 Recovering underlying distribution
Posteriorstacker uses the hierarchical Bayesian model described in Appendix A of Baronchelli et al. (2020) to infer the intrinsic distribution for desired parameters given posteriors for a sample of objects. We briefly describe the method here. The parent distribution of a value $n$ for all objects is assumed to be a Gaussian $N(n|\mu,\sigma)$ with unknown mean $\mu$ and standard deviation $\sigma$ (the hierarchical model’s hyperparameters); in this case $n$ is the braking index. As the parent distribution should hold for all objects, the hierarchical Bayesian modelling likelihood for the hyperparameters can be calculated as the product of the individual pulsar likelihoods marginalised over $n$
@@
where $\mathbf{D}$ refers to the full dataset of all pulsars. The posterior distributions on the braking index, $P(n|D_{i})$, for each pulsar, $i$, can be used as likelihoods in equation (7), i.e.,
@@
where $m_{i}$ is the number of samples and $n_{ij}$ are the posterior sample values of the braking index for the $i^{\rm th}$ pulsar, i.e., it is the expectation value/mean of the parent distribution evaluated at the posterior sample values. So, the likelihood that posteriorstacker evaluates over $\mu$ and $\sigma$ can be approximated by
@@
Posteriorstacker provides two model distributions: histogram (using a Dirichlet prior distribution) and Gaussian. The histogram distribution is more agnostic of the true underlying distribution, but it is therefore also less constraining. We use the Gaussian distribution as it provides better constraints and has fewer parameters to infer (i.e., it’s a “simpler” model). For the Gaussian model, posteriorstacker estimates the distributions for the mean and standard deviation for the braking index. The solid lines in Figures 2-4 come from generating Gaussians using the median values from these distributions, while the shaded bands come from the values at the 95th and 5th percentiles of the distributions. Posteriorstacker can provide an evaluation of the parent Gaussian distribution at different percentiles of the sampled hyperparameters, i.e., the median (50th percentile) distribution.
To summarise, for each analysis presented below, we perform the following steps: i) generate simulated TOAs for a set of pulsars, ii) for each pulsar, use the simulated TOAs to draw posterior samples on $f$, $\dot{f}$ and $n$ using a nested sampling algorithm, and iii) use the combined posterior samples on $n$ from all pulsars to infer posterior samples on the mean $\mu$ and standard deviation $\sigma$ of a parent Gaussian distribution for $n$.
@@ 2.4 Quantifying the results
The results obtained in this analysis are plots, which provide easy visual analysis of the degree to which recovering $n$ is successful. We also want to be able to quantify these results numerically. The equation
@@
is the analytical solution to integrating the Gaussian distribution from $a$ to $b$, where $\mu$ and $\sigma$ are the mean and standard deviation.
First, we decide on an arbitrary grid on the braking index. Then for each grid bin we evaluate equation (2.4) using the grid bin boundaries, $b_{\text{min}}$ and $b_{\text{max}}$, as the limits of integration for the all the equally weighted posterior sample values distributions. We then sum the probabilities within each bin. For each bin, this gives the probability, $\text{O}_{b_{\text{min}}-b_{\text{max}}}$, that a braking index randomly drawn from our distributions would be within that bin:
@@
To find the equivalent probability within an arbitrary range, in our case $[4,6]$, we can sum up the probabilities for the bins within that range. We can then also sum together the probabilities for all bins, $\text{O}_{\text{all}}$ (in our case, this is $[0,10]$) to obtain the odds ratio:
@@
In this analysis, we choose to use 100 bins and calculate the odds ratio within the range $[4,6]$, denoted as $\text{OR}_{4-6}$. Figure 1 visualises the results of this process for a run with an observation length of 20 years, an RMS noise value of $1\!\times\!10^{-5}$ ms using all 47 pulsars. For this run, $\text{OR}_{4-6}=1.00$, meaning there is a 50% chance of drawing a value of $n$ between 4 and 6 from the distributions.
Figure 1: A normalised histogram of the equally weighted posterior samples binned into 100 bins for an observation length of 20 years, an RMS noise value of $1\!\times\!10^{-5}$ ms using all 47 pulsars. Using equation (2.4), probability that $n$ lies in the bins between 4 and 6 (shaded in red) and the probability that it lies outside that region is calculated to find $\text{OR}_{4-6}$.
@@ 3 Analysis results
Following the methods described in Section 2, we repeat the analysis several times, changing the parameters of the pulsar TOA dataset to identify those with the greatest impact on the recoverability of the braking index using this method.
@@ 3.2 Observation length
Since $n$ depends on the frequency double derivative which changes very slowly over time, we expect more accurate and confident results with increased observation time as the change over that period will be larger. This is especially true as $\ddot{f}$ will have a cubic increase with time. Additionally, $\ddot{f}$ decorrelates with $\dot{f}$ over time, making it easier to measure. Figure 3 shows the results of five runs of varying observation lengths. In this case, the simulated TOAs are created for each analysis, with different specified lengths. The mean RMS noise is kept at $1\!\times\!10^{-5}$ ms and we include all 47 pulsars, then identical posterior analyses is performed.
@@ 4 Discussion
Of the factors considered, improvement of some are more achievable than others. Although some pulsars have been observed for 50 years, such as Vela, having a large sample with consistent observations and small timing residuals for such a time will not happen soon. For example, the NANOGrav dataset (Alam et al., 2020) used in this analysis has 12.5 years of observations for 47 pulsars, while the EPTA has 24 years for 6 pulsars (Chen et al., 2021) and the PPTA as observed 30 millisecond pulsars spanning up to 18 years (Reardon et al., 2023). The newest NANOGrav dataset, the 15 year dataset, has 67 pulsars (Agazie et al., 2023). These three groups are part of the International Pulsar Timing Array (IPTA). In the future, combined arrays such as this will allow for much larger populations of pulsars to be analysed for longer periods of time, thus increasing the likelihood that this analysis would observe the underlying distribution on the braking index.
@@ Appendix A Tables of parameters
Table 3: Pulsar values used in this analysis. Columns denoted by include values calculated from a simulated 20 year run with an RMS of $1\!\times\!10^{-5}$ ms assuming all pulsars have a braking index of exactly $n=5$.

### v2

@@ Abstract
The braking index, $n$, of a pulsar is a measure of its angular momentum loss and the value it takes corresponds to different spin-down mechanisms. For a pulsar spinning down due to gravitational wave emission from the principal mass quadrupole mode alone, the braking index would equal exactly 5. Unfortunately, for millisecond pulsars, it can be hard to measure observationally due to the extremely small second time derivative of the rotation frequency, $\ddot{f}$. This paper aims to examine whether it could be possible to extract the distribution of $n$ for a whole population of pulsars rather than measuring the values individually. We use simulated data with an injected $n=5$ signal for 47 millisecond pulsars and extract the distribution using hierarchical Bayesian inference methods. We find that while possible, observation times of over 20 years and RMS noise of the order of $10^{-5}$ ms are needed, which can be compared to the mean noise value of $3\!\times\!10^{-4}$ ms for the recent wideband 12.5-year NANOGrav sample (Alam et al., 2021a), which provided the pulsar timing data used in this paper.
facilities: NANOGrav (Alam et al., 2021a, b)software: enterprise (Ellis et al., 2020), enterprise_extensions (Taylor et al., 2021, 2024), enterprise_warp (Goncharov, 2021; Goncharov et al., 2024) posteriorstacker (Buchner, 2021), bilby (Ashton et al., 2019; Romero-Shaw et al., 2020; Ashton & Talbot, 2021), numpy (Harris et al., 2020), Astropy (Robitaille et al., 2013; Price-Whelan et al., 2018), Tempo2 (Hobbs et al., 2006; Edwards et al., 2006; Hobbs et al., 2009), matplotlib (Hunter, 2007),
@@ 1 Introduction
$\dot{f}_{\rm dip}=-\frac{8\pi^{2}}{3c^{3}}\frac{B^{2}}{I_{zz}}R^{6}f^{3}\sin^{2}{\alpha},$ (2)
which gives a braking index of 3 (e.g., Hamil et al., 2015). Here, $c$ is the speed of light, $B$ represents the magnitude of the dipolar magnetic field at the magnetic pole, $I_{zz}$ is the moment of inertia, $\alpha$ is the angle between the magnetic field axis and the rotation axis, and $R$ is the star’s radius at the magnetic pole.
@@
$\dot{f}_{\rm GW}=-\frac{512\pi^{4}}{5}\frac{G}{c^{5}}I_{zz}\varepsilon^{2}f^{5},$ (3)
@@ 2.1 Creating TOAs
A sample of $47$ pulsars from the wideband NANOGrav 12.5-year data set (Alam et al., 2021a, b) are used as example pulsars for this analysis. All pulsars in this sample are MSPs, meaning they have a rotation frequency of over 100 Hz. This puts the sample closer to the cut-off described in Woan et al. (2018) which is caused by $n=5$ processes than if it contained pulsars with lower frequencies. It may be noted that 30 of the pulsars are in binary systems; however, this information is not used in this study and they are treated as isolated pulsars.
@@
In this analysis, we allow our simulated TOAs to be created with a root mean square (RMS) noise value for white noise alone. For each pulsar, the noise value is taken from Table 5 in Alam et al. (2021a). In cases where the source has a separate red noise value, the value for white noise is taken only. These values are listed in Table 3. As the effect of timing noise on the accuracy of this method will be tested, we need to be able to vary the average noise of the sample to any desired value $\text{RMS}_{\text{desired}}$. Therefore, we scale the RMS noise of each individual pulsar using
@@
where RMSi is the RMS noise value for pulsar $i$ taken from Alam et al. (2021a), $\mu$ is the mean of the original dataset, and $\mu_{\text{desired}}$ is the mean the sample will be scaled to. This retains the realistic natural variation of RMS noise throughout the population.
@@ 2.3 Recovering underlying distribution
Posteriorstacker uses the hierarchical Bayesian model described in Appendix A of Baronchelli et al. (2020) to infer the intrinsic distribution for desired parameters given posteriors for a sample of objects. We briefly describe the method here. The parent distribution of a value $n$ for all objects is assumed to be a Gaussian $N(n|\mu,\sigma)$ with unknown mean $\mu$ and standard deviation $\sigma$ (the hierarchical model’s hyperparameters); in this case the objects are pulsars and $n$ is the braking index. As the parent distribution should hold for all pulsars, the hierarchical Bayesian modelling likelihood for the hyperparameters can be calculated as the product of the individual pulsar likelihoods marginalised over $n$
@@
where $\mathbf{D}$ refers to the full dataset of all 47 pulsars used in this analysis. The posterior distributions on the braking index, $P(n|D_{i})$, for each pulsar, $i$, can be used as likelihoods in equation (7), i.e.,
@@
where $m_{i}$ is the number of samples (usually on the order of 20,000 per pulsar) and $n_{ij}$ are the posterior sample values of the braking index for the $i^{\rm th}$ pulsar, i.e., it is the expectation value/mean of the parent distribution evaluated at the posterior sample values. So, the likelihood that posteriorstacker evaluates over $\mu$ and $\sigma$ can be approximated by
@@
Posteriorstacker provides two model distributions: histogram (using a Dirichlet prior distribution) and Gaussian. The histogram distribution is more agnostic of the true underlying distribution, but it is therefore also less constraining. We use the Gaussian distribution as it provides better constraints and has fewer parameters to infer (i.e., it’s a “simpler” model). For the Gaussian model, posteriorstacker estimates the distributions for the mean and standard deviation for the braking index. Posteriorstacker can provide an evaluation of the parent Gaussian distribution at different percentiles of the sampled hyperparameters, i.e., the median (50th percentile) distribution. The solid lines in Figures 2-4 come from generating Gaussians using the median values from these distributions, while the shaded bands come from the values at the 95th and 5th percentiles of the distributions.
To summarise Section 2 so far, for each analysis presented below, we perform the following steps: i) generate simulated TOAs for a set of pulsars, ii) for each pulsar, use the simulated TOAs to draw posterior samples on $f$, $\dot{f}$ and $n$ using a nested sampling algorithm, and iii) use the combined posterior samples on $n$ from all pulsars to infer posterior samples on the mean $\mu$ and standard deviation $\sigma$ of a parent Gaussian distribution for $n$.
@@ 2.4 Quantifying the results
The results obtained in this analysis are plots, which provide easy visual analysis of the degree to which recovering $n$ is successful. We also want to be able to quantify these results numerically by calculating the odds ratio that a randomly drawn braking index would be within a given range about the desired value. To achieve this, we begin with the following equation
@@
which is the analytical solution to integrating the Gaussian distribution from $a$ to $b$, where $\mu$ and $\sigma$ are the mean and standard deviation.
First, we decide on an arbitrary grid on the braking index. Then for each grid bin we evaluate equation (2.4) using the grid bin boundaries, $b_{\text{min}}$ and $b_{\text{max}}$, as the limits of integration for the all the equally weighted posterior sample values distributions. We then sum the probabilities within each bin. This gives the probability, $\text{O}_{b_{\text{min}}-b_{\text{max}}}$, that a braking index randomly drawn from our distributions would be within that bin:
@@
To find the equivalent probability within an arbitrary range, for example $[4,6]$, we can sum up the probabilities for the bins within that range. We can then also sum together the probabilities for all bins, $\text{O}_{\text{all}}$ (in our case, this is $[0,10]$) to obtain the odds ratio:
@@
In this analysis, we choose to use 100 bins and calculate the odds ratio within the range $[4,6]$, denoted as $\text{OR}_{4-6}$. While the choice is arbitrary, this range is chosen as it excludes the $n=3$ case where the spin-down is due to magnetic dipole radiation alone and $n=7$ case where it is due to GW emission via r-modes. Figure 1 visualises the results of this process for a run with an observation length of 20 years, an RMS noise value of $1\!\times\!10^{-5}$ ms using all 47 pulsars. For this run, $\text{OR}_{4-6}=1.00$, meaning there is a 50% chance of drawing a value of $n$ between 4 and 6 from the distributions.
Figure 1: A normalised histogram of the equally weighted posterior samples binned into 100 bins for an observation length of 20 years and an RMS noise value of $1\!\times\!10^{-5}$ ms using all 47 pulsars. Using equation (2.4), probability that $n$ lies in the bins between 4 and 6 (shaded in red) and the probability that it lies outside that region is calculated to find $\text{OR}_{4-6}$.
@@ 3 Analysis results
Following the methods described in Section 2, we repeat the analysis several times, changing the RMS noise, observation length, and number of pulsars to identify the parameters with the greatest impact on the recoverability of the braking index using this method. In the following sections, the “default” parameters will be an observation length of 20 years with a mean RMS noise of $1\times 10^{-5}$ ms using all 47 pulsars. The run using these parameters is depicted as the solid blue line in each plot.
@@ 3.2 Observation length
Since $n$ depends on the frequency double derivative which changes very slowly over time, we expect more accurate and confident results with increased observation time as the change over that period will be larger. This is especially true as $\ddot{f}$ will have a cubic increase with time. Additionally, $\ddot{f}$ decorrelates with $\dot{f}$ over time, making it easier to measure. Figure 3 shows the results of five runs of varying observation lengths. In this case, the simulated TOAs are created for each analysis, with different specified lengths going backwards in time. This could theoretically result in slightly better results due to higher pulsar $f$, $\dot{f}$, and $\ddot{f}$ values than in future measurements. However, this, and the location of the epoch within the observation length, was verified to make negligible difference. The mean RMS noise is kept at $1\!\times\!10^{-5}$ ms and we include all 47 pulsars, then identical posterior analyses is performed.
@@ 4 Discussion
Of the factors considered, improvement of some are more achievable than others. Although some pulsars have been observed for 50 years, such as Vela, having a large sample with consistent observations and small timing residuals for such a time will not happen soon. For example, the NANOGrav dataset (Alam et al., 2021b) used in this analysis has 12.5 years of observations for 47 pulsars, while the EPTA has 24 years for 6 pulsars (Chen et al., 2021) and the PPTA as observed 30 millisecond pulsars spanning up to 18 years (Reardon et al., 2023). The newest NANOGrav dataset, the 15 year dataset, has 67 pulsars (Agazie et al., 2023). These three groups are part of the International Pulsar Timing Array (IPTA). In the future, combined arrays such as this will allow for much larger populations of pulsars to be analysed for longer periods of time, thus increasing the likelihood that this analysis would observe the underlying distribution on the braking index.
@@ Appendix A Tables of parameters
Table 3: Pulsar values used in this analysis. Columns denoted by include values calculated from a simulated 20 year run with an average RMS of $1\!\times\!10^{-5}$ ms scaled using the listed RMS value and equation (5) and assuming all pulsars have a braking index of exactly $n=5$.